The probability that the man will be alive 25 years is 2/3 and the probability that his wife will be alive 25 years 2/3.find the probability that:

both will be alive


1
Expert's answer
2022-04-05T13:10:29-0400

Let,  M = the event that the man will be alive in 25 years, and

W = the event that his wife will be alive in 25 years.

We are said that "P(M) =\cfrac{2} {3} , \ P(W) =\cfrac{2} {3}."

Observe that, M and W are independent events , so

"P(M\cap W) =P(M) \cdot P(W) =\cfrac{2} {3}\cdot \cfrac{2} {3}=\cfrac{4} {9}."